Solve: 2sin³x = cos x ( to find the General Solution) 

Grade:11

1 Answers

Vikas TU
7 years ago
Squaring both sides:
we get,
2sin³x = cos x
4sin^6x = cos^2 x
4sin^6(x) +  sin^2(x) =  1
sin^2x = t (let)
4t^3 + t – 1 = 0     {  0